x=2y=4
the directions are rewrite each linear equation in slope intercept for. Name the sow and y-intercept ad sketch the graph.
i have no clue how to do this. im on home bound. which means im out of school for the rest of the year. my teachers send me work but i have no clue how to do it. please help! can you show your work too. so ill know how to do the other questions by myself?!

Answer by mananth(16949)

I presume the equation ius x+2y =4
write in the form y = mx +b
x+2y=4
-x on both sides
x-x+2y= 4-x
2y=4-x
/2
y=2-1/2 * x
y = -1/2 x +2
compare with y= mx+b. where m is the slope and b= y intercept.
m= -1/2 theslope. y intercept = b = 2
y==-1/2 x +2
when x=0, y=2 . ( 0,2)
when y=0 x=4 . (4,0)
Plot the points draw the line . you are done